OpenStudy (anonymous):

HELP PLEASE! WIN MEDAL! Which statement is true concerning the graph of two lines showing the given rates? Rate 1: 20 ft in 5 s Rate 2: 18 ft in 3 s OPTIONS- A. The graph with Rate 1 passes through (3, 12). B. The graph with Rate 2 passes through (3, 15). C. The graph with Rate 1 passes through (8, 2). D. The graph with Rate 2 passes through (12, 2).

OpenStudy (anonymous):

@Harsha19111999

OpenStudy (harsha19111999):

Could you expand the question or do you have any options? Because I didn't understand the question

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

A. The graph with Rate 1 passes through (3, 12). B. The graph with Rate 2 passes through (3, 15). C. The graph with Rate 1 passes through (8, 2). D. The graph with Rate 2 passes through (12, 2).

OpenStudy (anonymous):

The equation for the line is y = slope * x For line 1: The rate or slope is 20 ft for 5s i.e 4ft/s For line 2: The rate is 6ft/s Now write the equations and see for the options. Hope this helps.

OpenStudy (anonymous):

I Don't get it :/

OpenStudy (anonymous):

Ok So line 1 is y = 4*x and line 2 is y = 6*x Put (3,12) and see if this satisfies any equation.

OpenStudy (anonymous):

Thanks! Win a Medal!

OpenStudy (harsha19111999):

Both C and D satisfies

OpenStudy (harsha19111999):

If you draw a graph for both of the rates and substitute C and D values, you can see that it satisfies

OpenStudy (anonymous):

:/ But I can only pick 1 option

OpenStudy (anonymous):

D does not satisfy.

OpenStudy (harsha19111999):

How?

OpenStudy (anonymous):

If it were (2,12), then it satisfies line 2.

OpenStudy (harsha19111999):

No. Check it once again you will get 6y=x

OpenStudy (anonymous):

The eqn is y = 6x The point is (12,2) 2 = 6*12 is not true. D cannot be a possibility.

OpenStudy (harsha19111999):

y-y1=y2-y1/x2-x1(x-x1) is the formula i used

OpenStudy (anonymous):

But you don't have two points to use this formula..

OpenStudy (harsha19111999):

you can get 2 points 20 ft in 5 s is 4 ft in 1 sec